OpenStudy (anonymous):

Find a cubic function with the given zeros. √(2),-√(2),-2

OpenStudy (anonymous):

Expand the following and simplify:\[\left(x-\sqrt{2}\right)\left(x+\sqrt{2}\right)(x+2)=0\]\[x^3+2 x^2-2 x-4=0\]